### v3

# Infusion: internal diffusion for inpainting of dynamic textures and complex motion

Abstract.
Video inpainting is the task of filling a region in a video in a visually convincing manner. It is very challenging due to the high dimensionality of the data and the temporal consistency required for obtaining convincing results. Recently, diffusion models have shown impressive results in modeling complex data distributions, including images and videos. Such models remain nonetheless very expensive to train and to perform inference with, which strongly reduce their applicability to videos, and yields unreasonable computational loads. We show that in the case of video inpainting, thanks to the highly auto-similar nature of videos, the training data of a diffusion model can be restricted to the input video and still produce very satisfying results. This leads us to adopt an internal learning approach, which also allows us to greatly reduce the neural network size by about three orders of magnitude less than current diffusion models used for image inpainting. We also introduce a new method for efficient training and inference of diffusion models in the context of internal learning, by splitting the diffusion process into different learning intervals corresponding to different noise levels of the diffusion process. To the best of our knowledge, this is the first video inpainting method based purely on diffusion. Other methods require additional components such as optical flow estimation, which limits their performance in the case of dynamic textures and complex motions. We show qualitative and quantitative results, demonstrating that our method reaches state of the art performance in the case of dynamic textures and complex dynamic backgrounds.
inpainting, video, diffusion models, editing
submissionid: papers_1080ccs: Computing methodologies Image processing
Figure 1. Our method can inpaint videos with dynamic textures and complex motion using a diffusion-based approach (inpainted region in blue in the film reel). High quality results are obtained with a reasonable network size by learning on the input video only, and leveraging our proposed training strategy: interval training. Training and inference are mixed during the diffusion process, which is divided into intervals.
[…]
A video inpainting example of our method, with a person in front of a fireplace inpainted. Close-up view of the fire reveals that the completion is realistic and dynamic.
1. Introduction
Video inpainting, or video completion, aims to fill in an unknown region in a video, in a visually convincing manner. While in the case of images, very impressive results have been obtained using deep learning approaches (Yu et al., 2018; Saharia et al., 2022a), the extension of these methods to video has proven to be extremely difficult. This is due to the great technical challenges of videos, such as the large dimensionality of the data which leads to computationally expensive methods, the high sensitivity of the human eye to temporal inconsistencies, and the difficulty of inpainting complex motions. Many current state of the art methods combine deep learning with optical flow to deal with temporal consistency and motion. Unfortunately, optical flow breaks down in the presence of complex motions, for example in the case of dynamic (video) textures, which are poorly synthesized by such methods, as we show in this paper.
Concurrently, diffusion models have recently emerged as a powerful framework to produce samples of complex distributions, allowing the synthesis of realistic and diverse examples of images, with unprecedented quality (Ho et al., 2020). However, these networks are extremely large (up to 500 million parameters for image inpainting (Lugmayr et al., 2022; Saharia et al., 2022a)), leading to very long training and inference times. While several acceleration techniques have been proposed (Song et al., 2021; Rombach et al., 2022) in the case of images, they would nevertheless yield prohibitive training times for video inpainting.
In this paper, we propose a diffusion-based video inpainting method which has both a reasonable network size (around 500k parameters), and avoids relying on optical flow that can lead to poor reconstruction of complex motions. We show that a lightweight diffusion network, combined with an “internal learning” method, produces high-quality video inpainting results, especially in the difficult case of complex motion. Internal learning refers to restricting the training database to the input video, which makes sense, due to the highly auto-similar nature of videos (Wexler et al., 2007). In particular, we propose a novel training algorithm for diffusion models, referred to as “interval training”, which allows us to greatly reduce the size of the network, leading to feasible training and inference times (between 3-15 hours on a single GPU). Furthermore, to the best of our knowledge, our work yields the first purely diffusion-based video inpainting algorithm, we avoid using the aforementioned optical flow. Thus, in contrast to most of the other video inpainting algorithms, our method handles dynamic textures well and is able to produce diverse inpainting solutions. Finally, temporal coherence is ensured by 3D convolutions in the network.
We evaluate our method quantitatively and qualitatively on several databases, displaying its ability to inpaint complex motions such as dynamic textures and moving objects. We show that optical flow-based methods fail in the case of dynamic textures, due to the stochastic nature of such motion. Our method, on the other hand, is purely based on diffusion, which is designed precisely to sample complex distributions. This confirms that diffusion is a good choice for such stochastic motion components.
2. Related work
2.1. Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have shown great generative capabilities on image datasets. They have supplanted Generative Adversarial Networks (GAN) in quality and training stability (Dhariwal and Nichol, 2021). One major drawback, however, is the slow inference and training, which has lead to several attempts at acceleration up with various techniques (Rombach et al., 2022; Nichol and Dhariwal, 2021).
For image inpainting, Lugmayr et al. (2022) use an unconditional diffusion model, with projections and time-travels for better coherence. Saharia et al. (2022a) train a conditional model for diverse image-to-image problems.
However, the extension of diffusion models to videos is not straightforward, in particular due to their huge sizes and training times. Blattmann et al. (2023) finetune an image model for videos by introducing temporal attention layers. Another option is to decompose the video as multiple images via neural atlases (Kasten et al., 2021) which can then be handled by image-based diffusion models (Chai et al., 2023). Ceylan et al. (2023) use noise inversion and attention maps for consistent video editing.
For video generation, Ho et al. (2022b) have extended the image diffusion approach by using a temporal attention layer for sharing the information along the temporal axis, avoiding costly 3D convolutions. Other approaches have studied the conditional generation of videos via multiple networks for frame interpolation and super-resolution (Ho et al., 2022a; Singer et al., 2023). Harvey et al. (2022) implement a diffusion network for generation based on long temporal attention and frame interpolation. To reduce the computational complexity of handling full 3D videos, Yu et al. (2023) use multiple 2D projections. Mei and Patel (2023) have two models for frame content and frame motion. The very recent diffusion model Sora produces videos from text with unprecedented quality. While the complete architecture and training strategy is not known at the time of writing (Brooks et al., 2024), it is likely that training or even running such a model requires unreasonable computing facilities.
Balaji et al. (2023) propose to create an ensemble of diffusion networks specialised in different steps of the diffusion process. This approach shares some similarities with our training method (it was developed concurrently to our work), however a major difference is that theirs does not address the issue of network size, which is so crucial in our video inpainting problem.
Related to the internal learning strategy proposed in our work, it has also been shown recently that diffusion can be applied in a small data setting, for instance when only a single image is available for training. Kulikov et al. (2023) thus achieve single image generation using only one image. Similarly, Nikankin et al. (2023) show the application to image as well as video (single video generation / extrapolation). It is also possible to finetune a general diffusion model for a specific scene (Ruiz et al., 2023).
2.2. Video Inpainting
Classical approaches to video inpainting have long relied on available information already present in the video to achieve inpainting of the occluded region. They rely on the auto-similarity hypothesis, which is the foundation for patch-based methods in inpainting (Wexler et al., 2007; Newson et al., 2014), as well as other problems such as denoising (Buades et al., 2005) or image editing (Pritch et al., 2009).
Another class of approaches is based on optical flow. Indeed, in many situations, especially when the occluded region is observed at some point in time in the video, optical flow gives very useful information. Huang et al. (2016) jointly optimize the appearance and the optical flow, making sure that the output is temporally coherent.
More recently, Kim et al. (2019) have proposed a direct inpainting approach using a 3D neural network. Methods such as (Xu et al., 2019; Gao et al., 2020) first inpaint the optical flow and then propagate pixel values. These methods have been successful in handling non-rigid motions and long range dependencies. Zhang et al. (2022c) proposed to predict the optical flow using more frames, resulting in a more accurate optical flow and better preservation of motions. Li et al. (2022) integrate the optical flow prediction and propagation inside the training loop, optimizing the network parameters for video inpainting end-to-end.
Transformer-based architectures have also been used for video inpainting (Liu et al., 2021; Zhang et al., 2022a). Transformers can handle long range dependencies needed in video inpainting, however, their heavy computational cost can be prohibitive for the large dimensionality of videos and imply drastic approximations of the attention. ProPainter (Zhou et al., 2023a) improves the long range propagation of known regions with Transformers.
Internal learning for video inpainting has been proposed by Zhang et al. (2019). They learn a convolutional model on a single video similarly to Deep Image Prior (Ulyanov et al., 2018), and model the optical flow for better temporal coherency. Ouyang et al. (2021) tackle the same problem but propose a solution to avoid explicit modeling of the optical flow. Finally, Ren et al. (2022) use a Transformer with a discrete codebook finetuned on the target video, similar to a masked quantized autoencoder.
Figure 2. Our UNet architecture used for denoising the video in our diffusion approach is very simple, using mainly 3D convolutions.
[…]
Our UNet architecture uses 3D convolutions, downsampling and upsampling operations. The number of layers is low and is easy to illustrate.
3. Method
3.1. Background
Figure 3. We train a diffusion model on the input video only using interval training. In interval training, we learn on a subset of timesteps and the inference is done immediately after the training phase. Starting from $x_{T}$, the inpainting result is progressively generated with a perfectly adapted network.
We first recall the diffusion framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), based on discrete forward and backward processes. We adapt it for inpainting by modeling the conditional data distribution $p(\mathrm{x}\mid y)$ where $\mathrm{x}$ is a full video, and $y$ are the observations i.e. the known (unoccluded) region of the video. We represent an RGB video of $L$ frames of size $H\times W$ with a vector in $\mathbb{R}^{L\times H\times W\times 3}$. Starting from the unknown distribution of completed videos $q(\mathrm{x}_{0}\mid y)$ conditioned on $y$, we define a Markov chain $q(\mathrm{x}_{0}\mid y),...,q(\mathrm{x}_{T}\mid y)$ with the forward transition kernel for $t\in[1,T]$:
(1) $q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1},y)=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\bm{I}\right),$
with the variance schedule linearly increasing from $\beta_{1}=0.0001$ to $\beta_{T}=0.02$. Associated with the forward process is the reverse process defining the data distribution $p_{\theta}(\mathrm{x}_{t}\mid\mathrm{x}_{t+1},y)$ which is also a Gaussian of unknown but learnable mean and variance:
$p_{\theta}\left(\mathrm{x}_{t}\mid\mathrm{x}_{t+1},y\right)=\mathcal{N}(\mathrm{x}_{t};\mu_{\theta}(\mathrm{x}_{t+1},y,t),\sigma^{2}_{t}\bm{I}).$
We the classical approach of a neural network to predict the conditional mean $\mu_{\theta}(\mathrm{x}_{t+1},y,t)$ whose parameters $\theta$ are optimized to maximize a lower bound of the log-likelihood of the data. Following (Ho et al., 2020) we use a fixed variance schedule for the reverse process: $\sigma^{2}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$.
In the case of inpainting, we are interested in the conditional data distribution $p(x\mid y)$ where $y$ are our observations i.e. the known region of the video, we thus train our network with additional inputs. We minimize the L2 loss of the $x_{0}$-parameterization (Ho et al., 2020):
(2) $\mathbb{E}_{\mathrm{x}_{0},\epsilon,t,M}\lVert M\circ\left(\mathrm{x}_{0}-f_{\theta}(\tilde{\mathrm{x}}_{t},y,t,M)\right)\rVert_{2}^{2},$
where $M$ is a binary mask containing 00 for the known region and $1$ elsewhere, $y$ is the masked video i.e. $y=x_{0}\circ(1-M)$, and $\tilde{\mathrm{x}}_{t}$ is obtained by the forward diffusion of $\mathrm{x}_{0}$ by the following:
$\tilde{\mathrm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}\mathrm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{i=0}^{t}\alpha_{t}$, $\epsilon\sim\mathcal{N}(0,I)$. The observations and masks are introduced via concatenation in the input layer (similarly to (Saharia et al., 2022a)).
Unless specified, we use the default parameters of (Ho et al., 2020): variance schedule, sampling, time schedule. Therefore, our method could benefit from many of the recent improvements for each of these items.
3.2. Architecture
For temporal coherency, we extend the traditional UNet architecture to videos by using 3D convolutions. Another option is to rely on optical flow, but this can be inaccurate for complex situations. Another common approach for long-range dependencies is to use the attention mechanism (Vaswani et al., 2017), but it is extremely computationally heavy, and prohibitively so in the case of videos. Our methods avoids the use of this mechanism, yielding a network with reasonable computational requirements.
Since we are only training the model on a single video, we can greatly reduce the number of parameters in our network in comparison to standard diffusion models. The network has a total of 16 convolutional layers, distributed over 4 levels. Our 3D UNet downsamples only in the spatial dimensions to preserve as much as possible the temporal dimension, which is already much smaller than the others. We only use 32 output channels in each layer resulting in only 500k parameters which are to be compared to the 500M parameters of Palette (Saharia et al., 2022a) for image inpainting and the billions of parameters of other video diffusion models (Ho et al., 2022a). Our code can be found in the supplementary materials.
3.3. Single Video (Internal) Training
Our neural network is trained to inpaint a single video, using the video itself. This type of approach is referred to as internal learning. Fundamentally, it is based on the auto-similarity hypothesis, which has been successfully used in the case of patch-based methods (Newson et al., 2014) and neural networks via internal learning (Zhang et al., 2019; Ouyang et al., 2021). The main assumption here, explaining the success of such methods for video inpainting, is that videos contain roughly the same content in successive frames, with some motion. with some or all of the occluded content being revealed at some point in time. Thus, what might be considered as overfitting in a normal machine learning situation is in fact desirable here: we want to use the content of the video to inpaint itself. Put another way, we take advantage of both the learning and auto-similarity paradigms in our method. This kind of situation also arises in the problem of Single-Image Generation  (Shaham et al., 2019).
It is thus possible to efficiently inpaint a video without requiring a large external dataset. We assume that a binary test mask $M_{\text{test}}$ is provided. It is equal to 1 in the region to be inpainted, 0 elsewhere. Training is done by sampling consecutive frames from the video, and generating training masks which are to be inpainted / denoised by our neural network. We prevent data leakage from the test region by additional masking operations for the inputs:
$\begin{array}[]{lcl}\tilde{x}_{t}&=&\sqrt{\bar{\alpha}}_{t}(1-M_{\text{test}})\circ x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\\ y&=&(1-M_{\text{test}})\circ(1-M_{\text{train}})\circ x_{0}\\ M&=&1-(1-M_{\text{test}})\circ(1-M_{\text{train}})\\ \end{array}$
$M$ is thus the logical OR combination of $M_{\text{test}}$ and $M_{\text{train}}$. Similarly, we ignore some parts of the objective function:
$\mathcal{L}_{\theta}(x_{0})=\lVert M_{\text{train}}\circ(1-M_{\text{test}})\circ\left(x_{0}-f_{\theta}(\tilde{x}_{t},y,M,t)\right)\rVert^{2}$
3.4. Interval training
Training a diffusion network obviously involves training it for all timesteps $t$, from $1$ to $T$, since all these steps are needed during inference. However, it is clear that, depending on the timestep, the goal of the network, to estimate $x_{0}$ will be very different: if $t$ is large (pure noise), then the network should behave more like a generative model. If $t$ is small, then the network should act more like a denoiser, since the original video is only slightly noisy. This observation is crucial: indeed, in a standard diffusion model, one single network is supposed to achieve these very different goals. This is one reason why diffusion networks tend to be extremely large.
To circumvent this problem, we propose to use one lightweight network, trained on one subset of timesteps at a time. The training starts from the end of the diffusion process on the first interval. Once training is done on this interval, we use the model to perform the inference on the test data for all timesteps of the interval, up to the beginning of the next interval. The model is then trained on the next time interval. This is carried out until we have reached time step $t=0$. We can, of course, save the intermediate models if we wish, but this is not really necessary for video inpainting, since the model can only be used for the current video.
In practice, we define a time schedule of $N$ steps such that $\tau_{1}=1<\tau_{1}<...<\tau_{N}=T$. Our network learns successively on each interval $[\tau_{i},\tau_{i+1}]$, starting from the most noisy interval. After training on an interval, the inference is done for the same interval i.e. for all timesteps from $\tau_{i+1}$ to $\tau_{i}$. Training is detailed in Figure 3 and Alg. 1.
Algorithm 1 Training / inference algorithm for interval training
model $f_{\theta}$, intervals $\tau_{i=1..N}$, test data $y_{\text{test}}$
$x_{\text{test}}\leftarrow\mathcal{N}(0,I)$
for $i=N-1,...,1$ do
     for $j=1,..,K$ do $\triangleright$ Training for $K$ iterations
         $x_{j},y_{j},\epsilon_{j},M_{j}\leftarrow\text{NextBatch}(j)$
         $t\leftarrow\mathcal{U}(\tau_{i},\tau_{i+1})$
         Take gradient step with gradient:
         $\quad\nabla_{\theta}\lVert x_{j}-f_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{j}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{j},y_{j},t,M_{j})\rVert^{2}$
     end for
     for $t=\tau_{i+1},...,\tau_{i}$ do $\triangleright$ Inference
         $\mu_{\theta}\leftarrow f_{\theta}(x_{\text{test}},y_{\text{test}},t)$
         $x_{\text{test}}\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{t}I)$
     end for
end for
return $x_{\text{test}}$
Table 1. Inpainting metrics on dynamic texture datasets. Our methods perform the best in terms of LPIPS and SVFID. Interval training brings a noticeable improvement over the baseline approach to diffusion. Among the recent methods, only our method is able to produce diverse results.
Method Dynamic textures (Tesfaldet) Dynamic textures (DTDB)
LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑ Diversity ↑ LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑
Patch (Newson et al., 2014) 0.051 0.227 29.47 0.954 0.053 0.0180 0.055 36.52 0.988
ISVI (Zhang et al., 2022c) 0.062 0.500 27.41 0.945 0 0.0302 0.248 34.20 0.982
E2FGVI (Li et al., 2022) 0.046 0.288 31.58 0.961 0 0.0145 0.078 38.76 0.991
FGT (Zhang et al., 2022b) 0.046 0.202 30.88 0.956 0 0.0144 0.059 37.99 0.990
ProPainter (Zhou et al., 2023a) 0.046 0.261 31.77 0.960 0 0.0137 0.051 39.62 0.991
Infusion (no interval) 0.042 0.145 29.54 0.960 0.039 0.0150 0.043 37.09 0.989
Infusion (ours) 0.035 0.116 30.27 0.964 0.054 0.0137 0.036 37.20 0.989
This strategy has several advantages. First, we can use a lighter network as it specializes only on a small subset of timesteps. Second, we can avoid complicated weighting schemes for the loss function: in one interval, all timesteps have roughly the same weight and the weighting term can thus be set to $1$. This is not the case when one network is used and must perform well on all $T$ timesteps (Hang et al., 2023; Nichol and Dhariwal, 2021).
4. Experiments
4.1. Evaluation and masks
We compare our method with several state of the art methods. We include a classical patch-based approach (Newson et al., 2014) which is also internal and well adapted to the inpainting of dynamic textures, as well as more recent deep learning approaches: E2FGVI (Li et al., 2022), ISVI (Zhang et al., 2022c), FGT (Zhang et al., 2022b), and ProPainter (Zhou et al., 2023a).
We evaluate our method on complex videos featuring dynamic textures, large occlusions, and moving objects. These include, firstly, video sequences with a segmented object to be removed, from (Newson et al., 2014; Granados et al., 2012; Le et al., 2019), as well as some new sequences also with segmented objects. Secondly, we use dynamic texture databases: (Tesfaldet et al., 2018), featuring 59 textures, and a subset of DTDB (Hadji and Wildes, 2018) containing one random video from each appearance category. In these last two databases, we use synthetic masks, since there are no meaningful objects (people walking, etc) to be removed.
We evaluate our method qualitatively on the first set of sequences, with objects to be removed, for which no ground truth is available. We evaluate our method quantitatively on the second set of sequences, using various classical and perceptual metrics.
A database which is often used to evaluate generic video inpainting methods, not included in our evaluation, is the DAVIS database (Perazzi et al., 2016). This database was initially developed to benchmark video segmentation methods. It contains a large amount of videos with strong camera motion, for which the optical flow is of crucial importance. On this database, methods combining optical flow and deep learning are clearly superior to ours, which is not adapted to such situations.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
FGT ProPainter Infusion
Figure 4. In these scenes, our method satisfactorily inpaints dynamic textures preserving sharp details (top). It also better handles complex occlusions such as people crossing paths.
4.2. Training details
We resize the training data and masks to 432x240 for sequences larger than that. While our architecture is lightweight and may thus be used with many frames on a reasonable GPU, we only use 20 frames at the same time for training. Contrary to other approaches for video generation, we use all consecutive frames without dropping any of them.
For training, we generate masks using random strokes similarly to Kim et al. (2019). Training masks may overlap with the test masks, in which case no loss is actually computed in the intersection.
We use the Adam optimizer with an initial learning rate of 1e-4, and train for 200k total steps. For interval training, the 1000 timesteps are divided into 20 intervals of length 50. The same training budget is allocated to each interval.
4.3. Quantitative results
We quantitatively evaluate the methods using Peak Signal-to-Noise Ratio (PSNR), structure similarity measure (SSIM), Learned Perceptual Image Patch Similarity (Zhang et al., 2018) (LPIPS) and Single Video Fréchet Inception Distance (SVFID) (Wang et al., 2018; Gur et al., 2020). Finally, we also compute a diversity score following (Shaham et al., 2019). As explained above, the evaluation is conducted on the two databases of dynamic textures.
The first three metrics (PSNR, SSIM and LPIPS) are computed frame-by-frame and therefore do not reflect motion discontinuities. SVFID, on the other hand, takes spatio-temporal features into account. For computing the SVFID, we follow (Dorkenwald et al., 2021) and use the same model, trained on dynamic textures, for extracting the spatio-temporal features required for its computation.
Results are given in Table 1. We first observe that the LPIPS metric is significantly better with our method. In contrast, PSNR and SSIM are relatively good for all methods, the SSIM being slightly better for us and the PSNR better for the two deep-learning based methods (Zhou et al., 2023a)and (Li et al., 2022). A visual inspection of the videos provided with this paper nevertheless clearly shows that such methods struggle to reproduce spatial details of the textures. This can be explained by the fact that the PSNR is well-known to be poorly adapted to fine details (Blau and Michaeli, 2018; Saharia et al., 2022b). The LPIPS metric, on the other hand, is more faithful to human perception.
We also observe on Table 1 that the SVFID metric, the only truly video-based metric considered here, is significantly better for our approach. This shows that our method respects spatio-temporal statistics better than competing methods. In fact, a visual inspection of the resulting video (included in the supplementary materials) shows that for such content, deep-learning based competing approaches (ISVI, FGT, E2FGVI, and ProPainter) often produce results that are almost still or have completely wrong dynamics. This behavior is likely to be the result of a heavy use of optical flow to extend two-dimensional deep learning-based methods to videos. In addition, these methods have been trained on general purpose video databases, in which dynamic textures are often poorly represented. In contrast, the internal training strategy of our method and the fact that it is based on the diffusion framework with a purely convolutional network (no optical flow) is especially adapted to dynamic textures and high spatial and temporal frequency.
The diversity score shows a superiority of our approach and of the patch-based method from (Newson et al., 2014) with respect to other approaches. In fact, all considered deep-learning methods always produce the same result, yielding a null diversity score.
The diversity of our algorithm outcomes also enables us to perform the following experiment. For each texture of the Tesfaldet dataset, we produce 100 outcomes and average them to produce a new result, included in Table 2 as infusion (avg). This averaged result has a significantly better PSNR score than single outcomes. As expected, this improvement in PSNR comes at the cost of poorer fine details (Figure 10 and videos in the supplementary materials) and lower perceptual and diversity scores, more in line with the performance of the best methods among baselines E2FGVI, ISVI, FGT, and ProPainter. This is not surprising since these methods are not generative models but regressors. We believe that this simple experiment cleary higlights the strong limitation of the PSNR to evaluate results, in particular in presence of high frequency content.
Table 2. Inpainting metrics when averaging 100 samples or using a single one on the Tesfaldet dataset. PSNR is improved by averaging but the perceptual metrics are severely degraded.
[2] Method LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑ Diversity ↑ Infusion (avg) 0.056 0.231 30.97 0.962 0 Infusion 0.035 0.116 30.27 0.964 0.054
4.4. Object removal experiments
In this section, we consider the problem of object removal through several classical sequences used in previous works on video inpainting (Newson et al., 2014; Granados et al., 2012; Le et al., 2019) and a few new ones. We consider the same methods as before. For each sequence, a mask is given corresponding to an object (often a person) to be removed. As already mentioned, there is no ground truth for these sequences, so that quantitative evaluation is not performed. Nevertheless, such situations correspond to real applications and a more realistic foreground-background articulation than with synthetic masks. All results can be found in the supplementary materials. One can observe a much better respect of backgrounds such as dynamic textures or occluded moving objects with our approach than with other deep-learning based ones. This is briefly illustrated in Figure 4. On the first line, one sees that the unstable reflection of the building in the water is better reproduced by our methods than by FGT and ProPainter. The second and third rows display cases with occluded moving objects. In this case, results are blurry for ProPainter and blending artifacts are visible for FGT, while ours yields more satisfactory structures. The reader is strongly advised to look at the videos in the supplementary materials. These videos support the claim that a diffusion framework with appropriate internal learning is a good choice for stochastic and complex content.
4.5. Ablation study
This section demonstrates the importance of interval training. We compare the baseline training (i.e. training on all timesteps) and interval training. We train a neural network with the same number of parameters and for the same total number of training steps for both approaches. Two visual results can be seen in Figure 5. We observe that interval training clearly improves the reproduction of fine, textured details. More results can be seen on the videos in the supplementary materials. This observation is confirmed quantitatively in Table 1. Interval training significantly improves the quantitative results, especially for the perceptual metrics (LPIPS, SVFID) as well as for diversity.
We emphasize that in the baseline approach, results are very sensitive to the weighting scheme of the loss function (Nichol and Dhariwal, 2021; Hang et al., 2023). Our interval training approach has no such weighting, since each interval is trained separately, and uniform weighting on the interval is sufficient in all cases. Thus, our algorithm does not require time-consuming tuning of the loss weighting, which is a significant advantage.
Refer to caption Refer to caption
Refer to caption Refer to caption
Baseline training Interval training
Figure 5. Interval training clearly improves the visual results, especially the finer details. Associated videos are found in the supplementary materials.
4.6. Limitations and failure cases
Our method relies on a convolutional architecture which has a limited receptive field, especially in the temporal dimension. Our method may have trouble reproducing still background for long periods of time. This limitation is most visible for regions that are never seen in the video. We show an example in Figure 6 where the region inpainted changes from grass to water over a long occlusion (more than 30 frames apart) when it should stay stable.
Another limitation of our method is that each model is “single-use”, in the sense that it is adapted to a single video. This makes sense in the case of video inpainting, and greatly reduces network size and training time but it means that the model cannot be re-used, as it can be in other approaches. The optical flow is much more generalisable from one video to the next but, as we have seen, is limiting for elements such as dynamic textures. Thus, future research would involve creating a network which can be used on different videos.
Finally, our method is fully internal and thus relies on content from the video itself. Thus, it may have difficulty introducing completely new content in the video. Adding such completely new content, while maintaining a reasonable network size is a significant challenge for future research.
Refer to caption Refer to caption
Frame #65 Frame #99
Figure 6. When generating content, our method may not guarantee long term temporal consistency, especially for textures.
5. Conclusion
We have proposed a diffusion-based method for video-inpainting. To the best of our knowledge, this is the first purely diffusion-based approach to the video inpainting problem. This is made possible thanks to our novel training scheme and a lightweight network (around 500k parameters vs several billions for most video synthesis methods using diffusion). In particular, our training method leverages only the information contained in the current video, which is sufficient to achieve high quality inpainting, as confirmed by very satisfying results both quantitatively and qualitatively. We show a dramatic improvement for dynamic textures over current methods.
A future research direction would be to handle all situations, i.e. extreme camera motion and dynamic/complex scenes, within the same method. A second direction would be to accelerate training and inference times. To do so we can draw inspiration from Latent Diffusion Models (Rombach et al., 2022), distillation techniques (Salimans and Ho, 2022), or multi-scale diffusion models (Zhou et al., 2023b). Finally a major challenge is to extend the method to situations where the inpainted region should contain an element that does not resemble anything visible in the non-occluded part of the sequence.
Acknowledgments
This study has been carried out with partial financial support from the French Research Agency through the PostProdLEAP project (ANR-19-CE23-0027-01), and through the MISTIC project (ANR-19-CE40-005). Nicolas Cherel is supported by the Institut Mines Télécom, Fondation Mines-Télécom, and l’Institut Carnot TSN through the “Futur et Ruptures” grant.